OpenStudy (anonymous):

Compute the following derivative x/ln(x) I tried using the limit laws but I can't figure it out

OpenStudy (jdoe0001):

how about using the quotient rule? \(\bf \cfrac{x}{ln(x)}\implies \cfrac{\frac{d}{dx}[x]\cdot ln(x)-x\cdot\frac{d}{dx}[ln(x)] }{[ln(x)]^2}\)

OpenStudy (anonymous):

Ooooohh so I don't always have to apply the limit laws I can use quotient rule as well ?

OpenStudy (jdoe0001):

well... yes

OpenStudy (anonymous):

Okay thank you so much I got (1)*(lx(x))- (x/x) / (ln(x))^2 and after simplifying it makes more sense !

OpenStudy (jdoe0001):

yw